OpenStudy (anonymous):

differentiate the function: f(x)= x^4(1-(2/x+1))

OpenStudy (anonymous):

i just need help figuring out the f and g primes

OpenStudy (anonymous):

for the f prime i have 4x^3

OpenStudy (anonymous):

do you know tha chain rule as a generalization of diferenciating almost all kind of functions?

OpenStudy (anonymous):

but the g prime has me a little confused

OpenStudy (anonymous):

\[x^4(1-\frac{2}{x+1})\]?

OpenStudy (anonymous):

yes, like that satellite. also i'm familiar with the chain rule. i'm just having trouble with the g prime here

OpenStudy (anonymous):

my guess for the g prime was, ((-2/x+1)+1) = -2^-1x+1 + 1= 2x^-x

OpenStudy (anonymous):

the derivative of 1 is zero, and the derivative of \(\frac{1}{x}\) is \(-\frac{1}{x^2}\) so the derivative of \[1-\frac{2}{x+1}\] is \[\frac{2}{(x+1)^2}\]

OpenStudy (anonymous):

oh wow, i've never seen it like that. hmm

OpenStudy (anonymous):

you lost me on your guess, but it looks like you are trying to use the power rule while this rule always works, there are a couple derivatives you should memorize, because they come up again and again \[f(x)=\frac{1}{x}\] is a rather common function (take the reciprocal) and its derivative is \[f'(x)=-\frac{1}{x^2}\]

OpenStudy (anonymous):

you should also try to remember that the derivative of \(\sqrt{x}\) is \(\frac{1}{2\sqrt{x}}\) because the square root in another really really common function

OpenStudy (anonymous):

if you can remember these, and the chain rule, then you can see that the derivative of \[\frac{1}{f(x)}\] is \[-\frac{f'(x)}{f^2(x)}\] and the derivative of \[\sqrt{f(x)}\] is \[\frac{f'(x)}{2\sqrt{f(x)}}\] save you a ton of time on exams, when everyone else it trying to recall that \(\sqrt{x}=x^{\frac{1}{2}}\) etc

OpenStudy (anonymous):

i'll have to study those a bit more but i understand what you mean. thank you very much

OpenStudy (anonymous):

yw btw i really just mean "memorize" like knowing that \(7\times 8=56\) right off the top of your head good luck!